### v1

# ORFS-agent: Tool-Using Agents for Chip Design Optimization

Abstract
Machine learning has been widely used to optimize complex engineering workflows across numerous domains. In the context of integrated circuit design, modern flows (e.g., going from a register-transfer level netlist to physical layouts) involve extensive configuration via thousands of parameters, and small changes to these parameters can have large downstream impacts on desired outcomes - namely design performance, power, and area. Recent advances in Large Language Models (LLMs) offer new opportunities for learning and reasoning within such high-dimensional optimization tasks. In this work, we introduce ORFS-agent, an LLM-based iterative optimization agent that automates parameter tuning in an open-source hardware design flow. ORFS-agent adaptively explores parameter configurations, demonstrating clear improvements over standard Bayesian optimization approaches in terms of resource efficiency and final design metrics. Our empirical evaluations on two different technology nodes and a range of circuit benchmarks indicate that ORFS-agent can improve both routed wirelength and effective clock period by over 13%, all while using 40% fewer optimization iterations. Moreover, by following natural language objectives to trade off certain metrics for others, ORFS-agent demonstrates a flexible and interpretable framework for multi-objective optimization. Crucially, ORFS-agent is modular and model-agnostic, and can be plugged in to any frontier LLM without any further fine-tuning.
Index Terms:
Open-Source EDA, Physical Design, Autotuning, Large Language Model
I Introduction
Large Language Models (LLMs) have reshaped AI, excelling at natural-language generation, question answering, and zero-/few-shot learning [7]. Their emergent reasoning spans mathematical problem solving, code generation, and multi-step orchestration via agent frameworks [12]. LLMs integrate into agent systems [50], pairing API calls, predefined functions (“tool use”), and context windows to meet language-defined goals. These agents already serve fields from healthcare to voice-controlled computing.
Research on LLMs has seen major advances in neural architecture search [13] [32], combinatorial optimization [5] [37], and heuristic selection [6] [28], all relevant to chip design sub-tasks. Chip design relies on a complex EDA flow that turns high-level hardware descriptions into manufacturable layouts. Each stage exposes hundreds or thousands of hyperparameters (timing constraints, placement strategies, routing heuristics, etc.) whose joint settings strongly affect power, performance, and area (PPA). At advanced nodes, design complexity and the parameter space balloon further [1].
The Concrete Problem. We target LLM-driven black-box optimization across varied objectives, circuits, and technology nodes. In ML terms, this is context-based Bayesian optimization: LLM agents exploit metadata (e.g., log files) as heuristics. Traditional Bayesian optimization (BO) tunes EDA parameters well, yet pure black-box methods miss domain knowledge and need explicit objectives. LLMs, with broad “world knowledge,” use chain-of-thought to iteratively propose and refine parameters [62] [53].
LLM agents interface with chip-design tools, parse intermediate results, call domain APIs, and adjust flow parameters in a human-like loop. They also slot into existing heuristic frameworks (e.g., BO) to boost efficiency.
Distinction from other works: Fine-tuning vs. Model-Agnostic Approaches. Many recent efforts apply domain-specific fine-tuning. ChatEDA [53], ChipNeMo [33], and JARVIS [42] curate EDA-centric corpora, retrain or instruction- tune a base model, and then self-host the resulting checkpoint; this entails recurring data-collection, training, and maintenance costs whenever a stronger foundation model is released. ChatEDA, for instance, fine-tunes an open-source (or API-tunable) model on Q-A examples. Each new SOTA model demands another fine-tune: closed models (e.g., Claude) seldom allow this, and open models lag the frontier. Fine-tuning plus lifecycle management and hosting may cost USD \$10k, and the maintainer pays for inference.
Alternative approaches avoid retraining, and instead wrap or retrieve around frozen SOTA models. EDA-Copilot [4], Ask-EDA [47], and ORAssistant [25] combine retrieval-augmented generation with off-the-shelf chat models to answer tool-usage questions, while surveys (LLM4EDA [62], “Future or Mirage?” [60], and the taxonomy in [45]) highlight the growing appeal of lightweight, model-agnostic agents.
Our agent follows this latter paradigm: it wraps whatever SOTA model is available, with zero extra training cost; our full experiments can run on Grok-Mini for less than USD \$10. Further, while previous agents such as ChatEDA build general Q-A support for engineers, our agent focuses solely on optimizing flow parameters.
Our Contributions. We introduce ORFS-agent, an LLM-driven autotuner for an open-source chip-design flow that tops Bayesian optimization (BO) with fewer flow runs. Our main contributions are:
  • We embed ORFS-agent in OpenROAD-flow-scripts (ORFS)[70]. It schedules parallel ORFS runs, reads intermediate metrics (wirelength, timing, power), and iteratively refines parameters to optimize post-route results. Leveraging ORFS’s METRICS2.1 logging[22], the agent skips bad configurations and gives robust, efficient chip design optimization. Modularity allows instantly switching models.
  • ORFS-agent obeys natural-language instructions to weight or constrain design metrics (e.g., favor timing over wirelength) and taps existing BO libraries inside its tool-use loop for extra lift.
  • On ASAP7 and SKY130HD (three circuits each), ORFS-agent trims ORFS calls and improves wirelength and effective clock period by $\geq 13\%$ over baseline, beating the BO-based ORFS AutoTuner [22].
Figure 1: ORFS-agent integrated with OpenROAD-flow-scripts (ORFS); metrics gathered via METRICS2.1 [22].
In the following, Section II reviews BO and recent LLM-EDA work. Section III gives details of ORFS-agent—prompts, data, iterations. Section IV gives experiments. Section V concludes and sketches future work. Anonymized code is open-sourced for inspection at [71].
II Related Work
We organize our discussion of related work into three subsections, covering Open-Source EDA, Bayesian Optimization, and LLMs for Optimization in EDA in that order. Additional discussion of EDA flow and OpenROAD is given in Appendix A.
II-A Open-Source EDA and OpenROAD as an ML Testbed
We target LLM/ML EDA research based on the permissively open-sourced OpenROAD[3], where code, logs and data can be made public to enable reproducible training and evaluation. Tcl/Python APIs expose every flow stage, for capture of live metrics and debug – vital for ML agents. With its RTL-to-GDSII scope, transparency and active community, OpenROAD provides an excellent testbed. ORFS wraps OpenROAD in editable JSON/Tcl for large-scale experiments[70]. Other tools build on OpenROAD, e.g., DREAMPlace yields 30×30$\times$30 × faster placement than OpenROAD’s RePlAce via CUDA acceleration [30].
II-B Bayesian Optimization in Flow Tuning
Bayesian optimization (BO) is a widely used approach for parameter autotuning in EDA flows [22] including in commercial tools such as Synopsys DSO.ai [68] and Cadence Cerebrus [65]. It enjoys efficiency in terms of samples and solution space search, but faces drawbacks:
  • Contextual and Knowledge Limitations: BO frameworks typically do not incorporate domain knowledge unless manually encoded (e.g., at PDK level). Further, they do not adapt from past results.
  • Explicit Objective Functions: BO frameworks require explicitly defined objective functions, possibly missing complex or subjective criteria. This can hinder effectively balancing PPA in chip design.
  • Scalability Challenges: High-dimensional spaces and large designs can combinatorially overwhelm traditional BO approaches [57].
We seek flexible, context-aware and data-driven approaches to parameter tuning. Large Language Models offer a promising solution.
II-C Large Language Models (LLMs) in EDA Optimization
LLMs such as GPT [7] absorb web- and paper-scale corpora, giving them broad technical knowledge. They parse design logs, infer context, and suggest EDA optimizations [51] [62] [53].
Contextual knowledge and adaptivity. Unlike black-box Bayesian optimization (BO), LLMs ingest circuit data, retrieve literature, and adjust strategies on the fly. They read partial metrics (e.g., post- CTS timing) and learn user preferences without explicit objectives [40].
Recent progress. Advisor systems LLM4EDA and ChatEDA guide placement and timing fixes [62] [53]. Startups Silogy, Silimate, Atopile, Diode Computers and ChipAgents extend AI across design tasks. New methods such as LLAMBO [31, 69], forum posts on r/MachineLearning, and the LAD Symposium reinforce the trend. Multi-fidelity BO and meta-surrogates [14, 26, 46] improve sampling efficiency yet still treat flows as opaque; by contrast, LLMs promise context-aware tuning. Large-scale industrial adoption of LLMs for EDA can be seen, e.g., NVIDIA’s MARCO framework shows graph-based multi-agent task solving for design flows[66]. Open problems in LLM-based hardware verification are mapped in [54].
II-D Working with LLMs and Agent Frameworks
LLMs need scaffolding beyond raw prompts. Key techniques are:
  • Prompt engineering crafts system/user messages for clarity.
  • Chain-of-thought reasoning [51] elicits stepwise logic.
  • Retrieval-augmented generation [29] injects external documents.
  • Tool-use / agent frameworks [44, 15, 43] let an LLM call shells or solvers, parse error logs [20], and orchestrate full flows [27].
These methods enable adaptive, metric-driven tuning.
LLM agents for optimization. Agents excel at ill-specified goals [58], prompt search [16], hyperparameter tuning [34], autonomous research [19], and benchmark challenges [18]. EE Times has forecast the advent of collaborating AI agents for chip design by 2025 [11].
Coupling with Bayesian optimization. LLM-based agents can parse PPAC metrics, narrow BO search spaces, prune bad configs, or switch heuristics. Blending BO with symbolic reasoning [41] [49] yields targeted, adaptive flows. ORFS-agent builds on this foundation.
More modern advances. OPRO treats the model itself as the optimizer [58]. LLAMBO warm-starts BO with zero-shot hints [31]. ADO-LLM fuses analog priors with BO loops [59]. ChatEDA and the newer EDAid system wire agents straight into complete RTL-to-GDSII flows [53] [55]. RL efforts such as AutoDMP and Google’s graph-placement policy show tractability of end-to-end layout [2] [10] [38].
Tool-centric agent stacks. Browser agents such as WebArena surf, click, and scrape constraints for downstream solvers [61]. Terminal aides—ShellGPT and Devin—run make, git, and yosys on command, closing the loop between code and silicon [73] [67]. API routers like ToolLLM, Gorilla, and OpenAgents map natural-language tasks onto thousands of REST calls and plugins [44] [43] [56]. AutoGPT-style self-looping controllers review results and re-issue actions [48], while the AI-Scientist pipeline shows how the same pattern can automate hypothesis, experiment, and paper generation [74]. At code level, AlphaCode 2 beats 85% of Codeforces competitors [64]; RL-trained VeriSeek and retrieval-guided logic synthesis agents trim Verilog area-delay in minutes [35] [8]. Together, browser, tool, and terminal hooks turn LLMs into end-to-end optimization engines.
III Proposed Methodology
We now introduce the integration framework of our ORFS-agent with the OpenROAD-flow-scripts (ORFS). We then present an explanation of a single iteration of the ORFS-agent loop.
III-A Overview of ORFS-agent
ORFS-agent is an end-to-end agent framework designed to optimize OpenROAD-based chip design flows according to user-specified objectives and constraints. ORFS-agent leverages an LLM capable of performing internal text-based reasoning and invoking external tools (function-calling). Our approach supports parallel runs of ORFS, partial metric gathering and various optimization strategies. Figure 1 shows the overall framework of ORFS-agent.
The inputs to ORFS-agent include process design kits (PDKs) for the target technology node, target circuits described in Verilog files, and user-defined prompts. PDKs are sets of files that provide process-related information needed for chip implementation. In this work, we use two technology nodes: SKY130HD [72] and ASAP7 [63]. The prompts define the objective function and constraints in natural language. The objective function in this work is formulated as
$F=\sum_{i=1}^{n}\alpha_{i}f_{i},$
where each $f_{i}$ is a design metric (e.g., routed wirelength, worst negative slack) and each $\alpha_{i}$ is a nonnegative real weight. This formulation allows optimization of either a single metric or a weighted combination of multiple metrics. Then, ORFS-agent iteratively modifies the configuration file (config.mk) for ORFS and SDC files to optimize the specified objectives. The ORFS configuration file defines adjustable parameters (knobs) for placement, clock tree synthesis and routing. The SDC file specifies timing constraints for the target design, such as clock period. The outputs of ORFS-agent are the optimized configuration file for ORFS and the optimized SDC file; together these optimize the user-specified objectives under given constraints.
ORFS-agent relies on an LLM (e.g., GPT, Claude, Llama) that can: (i) read and modify files and logs, (ii) optionally invoke external tools via function calling, i.e., specify argument values for pre-existing functions (e.g., Python functions), and (iii) propose new parameter values to improve the objective under given constraints. There is a key distinction between direct LLM use and the use of LLM-based agents with function-calling capabilities. In the function-calling paradigm, the LLM is given a goal (e.g., “write a function to compute a matrix eigendecomposition”) along with a list of tools, which are black-box functions the LLM can invoke to facilitate task completion. For example, a tool named ${\mathbf{ISDIAGONALIZABLE}}$ determines whether a matrix $M$ is diagonalizable and returns a Boolean value. The tools used in ORFS-agent fall into the following three categories:
  • INSPECT tools allow the agent to inspect the data collected so far without context-level manipulation.
  • OPTIMIZE tools allow the agent to model and optimize the data by a suitable model such as a Gaussian Process.
  • AGGLOM tools allow the agent to sub-select from possible hyperparameters obtained from an OPTIMIZE tool.
In the basic function-calling agent loop, at iteration $t$ with context $C_{t}$, the agent performs the following up to K times:
  • Observe: Load the ORFS configuration file and SDC file, execute ORFS, and examine the results of ORFS;
  • Query: Call tools from the provided list (up to $M$ times) to gather information and update the context to $C_{t+1}$; and
  • Alter: Modify ORFS config/SDC files based on $C_{t+1}$.
By iterating through these steps, the function-calling agent incrementally refines its solutions by using external tools.
III-B Overall Iteration Structure with Example and Tool Usage
We describe one iteration of the ORFS-agent loop, which executes K parallel runs of OpenROAD-flow-scripts, with K being provided by the user. As shown in Figure 2, each iteration consists of the following steps: RUN, READ, COLLATE, INSPECT, OPTIMIZE, AGGLOMERATE (optional) and ALTER.
We maintain a GLOBALCONTEXT that collects relevant parameters and partial metrics across steps. The initialization of this GLOBALCONTEXT at the first iteration is also the LLM prompt; this initialization, along with procedures for update between iterations for GLOBALCONTEXT, is separately discussed in Appendix E. Let K denote the number of parallel runs per iteration, and let TIMEOUT specify the maximum runtime for each batch. Below, we show an example scenario with $\texttt{K}=25$, $\texttt{TIMEOUT}=30$ and parameters (Core Util, Clock Period) can be integrated, along with possible function-calling tools. Let us consider the $j$-th iteration. Assume that we target the routed wirelength and take the CTS wirelength as a surrogate (recalling that the CTS stage occurs earlier and thus will likely not time out).
Figure 2: Overall ORFS-agent flowchart.
Stage 1: RUN K Parallel Jobs. Launch K ORFS runs, each with distinct parameters, until the TIMEOUT is reached.
  • Example: For $\texttt{K}=25$ and $\texttt{TIMEOUT}=30$ minutes, we launch 25 parallel runs, each exploring a different (Core Util, Clock Period) setting. All will run for at most $30$ minutes, producing log files.
  • Tool use: This set of runs might be derived from a previous iteration where OPTIMIZE or AGGLOM tools suggest new parameter sets.
  • No tool use: If tool use is disabled, these parameter files arise from simple in-context operations.
Stage 2: READ. Gather JSON logs from the files generated by all RUN steps up to this iteration, which may contain partial surrogate metrics, e.g., from CTS stages as well as ground truth metrics from finished runs.
  • Example: The agent collects $25\times j$ JSON files, with the routed wirelength (ground truth) or the earlier, but rough CTS wirelength (surrogate).
  • No tools, hence identical with or without tool use.
Stage 3: COLLATE. Merge the logs into a single dataset, storing outcomes for all K ×$\times j$ runs (e.g., a table or array).
  • Example: Build a unified array with columns (Core Util, Clock Period, CTS wirelength, routed wirelength). The routed wirelength can be absent. If the CTS wirelength is absent, skip the row.
  • No tools, hence identical with or without tool use.
Stage 4: INSPECT. Analyze the collated data.
  • Example: For $\leq 25\times j$ ($\leq$ from skips) instances, find how Clock Period varies against the wirelength (actual target) and CTS wirelength (surrogate).
  • Tool use: The agent calls pre-defined INSPECT tools to examine distributions, which allows sidestepping inspecting the values in context. For example, $\mathbf{InspectDistribution(ARR,X,Y,Y_{surrogate})}$ outputs the distributional relation between an input $\mathbf{X}$ and targets $\mathbf{Y}$ with surrogate targets $\mathbf{Y_{surrogate}}$. Here we have Clock Period as $\mathbf{X}$, CTS and routed wirelength as $\mathbf{Y_{surrogate},Y}$ and $\mathbf{ARR}$ from the COLLATE step.
  • No tool use: The agent reasons purely in context using its chain of thought by inspecting the data.
Stage 5: OPTIMIZE. Based on INSPECT, propose new parameter sets for the next iteration.
  • Example usage: The agent calls a OPTIMIZE tool, e.g., a Gaussian Process ($\mathbf{GP}$) with the COLLATE-level data, to get $200$ values of candidate Clock Period, Core Util pairs, using a Matern kernel, and Expected Improvement as search strategy.
  • Tool use: In general, we may have candidate points from $\mathbf{GP}(\mathbf{ARR},\mathbf{args},\dots)$ or any other OPTIMIZE tool, obtaining many potential parameter sets. $\mathbf{ARR}$ arises from collate, and the agent decides on which OPTIMIZE tool to use, i.e., whether to use a Gaussian Process at all, as well as the choice of $\mathbf{args}$ depending on what it finds at the INSPECT phase. This may include choices over the kernel used, selection algorithm, and so on.
  • No tool use: The agent solely uses its own context to directly suggest a set of parameters.
Stage 6: AGGLOMERATE (optional). If more than K candidates are generated in the OPTIMIZE phase, then reduce to exactly K, often by maximizing diversity or coverage.
  • Example: Since the $\mathbf{GP}$ produces $200$ new (Core Util, Clock Period) pairs; we reduce to 25 via selecting through a valid AGGLOM tool, $\mathbf{EntropySelect(ARR,quality,points)}$. This function returns $\mathbf{points}$, i.e., $25$ pairs of (Core Util, Clock Period) that it thinks are optimal as a set w.r.t. the provided $\mathbf{quality}$ metric (wirelength improvement) on the same $\mathbf{ARR}$ from COLLATE.
  • Tool use: The agent may use any number of AGGLOM tool choices from the LLM’s decisions along with parameters to call them with. The final output is of shape K $\times$ the number of parameters optimized.
  • No tool use: Here, the OPTIMIZE case uses structured outputs to yield exactly K sets of parameter choices in context directly, and this step is skipped.
Stage 7: ALTER. Update ORFS configuration files with the newly selected parameter sets for the next iteration.
  • Example: Generate 25 updated config/SDC files (or commands), each specifying one of the chosen (Core Util, Clock Period) combinations.
  • No tools, hence identical with or without tool use.
This process repeats until the stopping criteria—generally, total serial iteration count—are met. The outputs of stages READ to AGGLOMERATE enter the context window of the ORFS-agent in each iteration, but not necessarily into GLOBALCONTEXT. Context management and tool-level details appear in Appendix E.
IV Experiments and Results
In our experiments, we use six design instances comprising three circuits (IBEX, AES and JPEG) in two technology nodes (SKY130HD and ASAP7). All experiments are performed on a Google Cloud Platform (GCP) virtual machine with 112 virtual CPUs (C2D AMD Milan) and 220GB RAM. The flow is based on the ce8d36a commit of the OpenROAD-flow-scripts repository, along with associated versions of OpenROAD, KLayout, Yosys, and other dependencies. [^1] To confirm reproducibility, we transfer the DEF files generated during the place-and-route process to another server and verify that the results are consistent across environments.
Although the timeout to run depends on the circuit, each “iteration” is consistently defined as 25 parallel runs, with the exception of the JPEG circuits, which are run with $12$ parallel runs each. Each of these parallel runs uses four vCPUs and 4 GB of RAM. These parallel runs write to different intermediate folders for logs and results, running Yosys and other steps in parallel without reusing intermediate steps. At the end of each iteration, a clean operation moves the logs and folders to a separate directory to avoid contamination.
Our experiments cover two settings. The first, “4-variable” setting involves optimizing over four parameters: Core Utilization, TNS End Percent, LB Add On Place Density, and Clock Period. The first three are in the configuration files of the respective circuits; Clock Period is in the .sdc file. The second, “12-variable” setting has eight additional tunable parameters that guide the physical design flow more comprehensively. We distinguish three experimental cases:
  • 4-variable, no tool use, where the LLM optimizes over four variables solely through prompt-tuning and in-context learning;
  • 4-variable with tool use, where Bayesian optimization enhances performance; and
  • 12-variable with tool use, which incorporates all 12 variables along with the optimization tools.
Our experimental choices reflect the LLM requiring more complex tools to deal with higher-dimensional optimization problems. All of our experiments involve either $375$ (4-variable setting) or $600$ iterations (12-variable setting). The ORFS-agent parameters are configured as follows: (i) the model used is claude-3-5-sonnet-20241022; [^2] We do not observe any performance gains from the o1-class of models. (ii) temperature is set to $0.1$; (iii) nucleus and top-K sampling are disabled; and (iv) to always output $25\times 4/12$ parameters (the number of parameters the LLM is optimizing over) in the specified format, we utilize JSON Mode (Structured Outputs API). [^3] In terms of API costs, all of our experiments, including all failed runs, cost USD \$48. Claude Sonnet 3.5 has since been overtaken by models which are simultaneously better and cheaper by 5-10$\times$ per token. As noted above, approaches such as ChatEDA require fine-tuning, which costs significantly more (by at least 10-20$\times$) and requires inference costs for serving the custom model. Appendix B provides additional experimental details.
Tech Circuit WL Optimization ECP Optimization Co-Optimization
WL ECP* ECP WL* WL** ECP**
OR-AutoTuner Performance: 375 Iterations, 4 Parameters
SKY130HD IBEX 632918 14.52 10.17 811265 802786 10.49
AES 514342 7.28 4.08 639876 564188 4.43
JPEG 1263740 8.25 6.77 1421110 1304829 7.24
ASAP7 IBEX 104622 1510 1252 109346 107821 1264
AES 68943 1634 431 78152 72303 446
JPEG 269721 1172 950 284539 276453 978
Average 1.00 1.00 1.00 1.00 1.00 1.00
OR-AutoTuner Performance: 1000 Iterations, 12 Parameters
SKY130HD IBEX 630598 14.08 10.09 767451 750769 10.10
AES 423730 5.02 3.81 563264 495781 3.89
JPEG 1070038 7.77 6.57 1295066 1089166 7.29
ASAP7 IBEX 96086 3244 1188 103081 100837 1192
AES 66934 2220 430 77067 71086 443
JPEG 260683 1130 882 279621 273649 885
Average 0.92 1.18 0.96 0.94 0.93 0.95
TABLE I: OR-AutoTuner results. Top block: 375 iterations with 4 parameters. Bottom block: 1000 iterations with 12 parameters. Columns are grouped by optimization: WL-only (WL, ECP*), ECP-only (ECP, WL*), and co-optimization (WL**, ECP**). Bold font indicates superior results.
Tech Circuit WL Optimization ECP Optimization Co-Optimization
WL ECP* ECP WL* WL** ECP**
375 Iterations, 4 Parameters without Tool Use
SKY130HD IBEX 732793 11.47 10.82 815984 792947 11.00
AES 531721 5.33 4.07 651219 637490 4.08
JPEG 1257231 7.28 6.69 1371607 1344380 6.77
ASAP7 IBEX 106044 1324 1281 113854 114576 1287
AES 70955 446 430 72427 71644 440
JPEG 277654 1084 1016 289031 286142 1069
Average 1.04 0.75 1.02 1.00 1.04 1.00
375 Iterations, 4 Parameters with Tool Use
SKY130HD IBEX 726874 10.91 10.74 752566 736275 10.84
AES 526782 5.12 4.02 601492 572966 4.06
JPEG 1268544 7.08 6.64 132785 1302494 6.74
ASAP7 IBEX 101788 1298 1242 112788 106854 1254
AES 70288 448 435 72144 71422 442
JPEG 270184 968 952 283544 273126 958
Average 1.03 0.71 1.00 0.82 0.98 0.97
600 Iterations, 12 Parameters with Tool Use
SKY130HD IBEX 657091 12.25 10.42 721099 706087 11.06
AES 483426 4.82 3.96 539878 500296 4.06
JPEG 1237829 6.94 6.54 1308171 1254030 6.62
ASAP7 IBEX 97305 1278 1248 105562 102205 1260
AES 69824 456 432 74582 71084 438
JPEG 270963 905 878 275944 271822 882
Average 0.98 0.71 0.98 0.92 0.94 0.96
TABLE II: ORFS-agent results for various tuning settings. Top: 375 iterations, 4 parameters (no tool); Middle: 375 iterations, 4 parameters (with tool); Bottom: 600 iterations, 12 parameters (with tool). Each row reports the figures of merit under WL-only, ECP-only, and Co-optimization objectives. Bold font indicates superior results.
[…]
  • $\mathbf{ECP}^{*}$ (resp. $\mathbf{WL}^{*}$) denotes the value of $\mathbf{ECP}$ (resp. $\mathbf{WL}$) when $\mathbf{WL}$ (resp. $\mathbf{ECP}$) is optimized as the objective.
  • $\mathbf{ECP}^{**}$ and $\mathbf{WL}^{**}$ represent the values obtained when both metrics are jointly optimized in a multi-objective setting.
We use OR-AutoTuner as the baseline. We run OR-AutoTuner for $1000$ iterations with $12$ parameters to optimize, and for $375$ iterations with $4$ parameters to optimize. Results are shown in Table I.
In the following, Section IV-A presents results for single-objective optimization. Then, Section IV-B gives results for multi-objective optimization. We then demonstrate ORFS-agent’s capability to perform constrained optimization guided by natural language commands. Appendix C provides additional ablation studies for the ASAP7 IBEX design. Studies on data leakage, model confidence and robustness are presented in Appendix D.
IV-A Single-objective Optimization
In the single-objective optimization scenario (e.g., minimizing routed wirelength $WL$), the LLM agent receives an additional natural language instruction appended to its prompt. Specifically, the instruction is:
There is only a single objective to optimize in this problem. You observe that the value inside the final JSON is $\mathbf{WL}$, while the corresponding baseline value is $\mathbf{WL_{\alpha}}$. Your effective loss to minimize should be $\mathbf{\frac{WL}{WL_{\alpha}}}$.”
where $\mathbf{WL_{\alpha}}$ denotes the routed wirelength obtained using the default configuration of OpenROAD-flow-scripts and serves as a fixed baseline for normalization. [^6] The baseline results obtained using the default configuration of OpenROAD-flow-scripts are shown in Appendix IV-C.
However, due to runtime timeouts, typically occurring during the detailed routing stage, the routed wirelength $\mathbf{WL}$ may not always be available. In contrast, the estimated wirelength at the end of CTS $\mathbf{WL^{\prime}}$ almost always exists. If $\mathbf{WL}$ is absent but $\mathbf{WL^{\prime}}$ exists, we include the following instruction:
You do not observe the correct value $\mathbf{WL}$, but you observe a strong surrogate $\mathbf{WL^{\prime}}$, which you should use as a signal. Your effective loss should be $\mathbf{\frac{WL^{\prime}}{WL^{\prime}_{\alpha}}}$.
where $\mathbf{WL^{\prime}_{\alpha}}$ denotes post-CTS wirelength obtained under the same default configuration. This surrogate-based guidance enables the LLM to continue learning from partial information while maintaining alignment with the optimization objective.
The experimental results are shown in Table II. Before discussing these results, we would like to highlight that our results are obtained with the Claude 3.5 class of models.
With respect to Table II, key observations are as follows.
  • Comparison between default ORFS and ORFS-agent: ORFS-agent (600 iterations, 12 parameters with tool use) achieves an average improvement of $13.2\%$ in wirelength (or ECP) when configured for wirelength (or ECP) optimization.
  • Comparison between OR-AutoTuner and ORFS-agent: Despite using 40% fewer iterations, ORFS-agent (600 iterations, 12 parameters, with tool use) achieves better results than OR-AutoTuner on $12$ out of $24$ metrics measured for single-objective optimization, thereby exactly drawing on cases of superiority.
  • Impact of tool use (375 iterations and 4 parameters): ORFS-agent with tool use dominates ORFS-agent without tool use in 5 out of 6 cases, for both wirelength and effective clock period optimization.
  • We notice that ORFS-agent does not significantly worsen the metric not being optimized. For instance, when optimizing wirelength on ASAP7 IBEX, the ECP obtained by ORFS-agent (600 iterations, 12 parameters with tool use) is $40\%$ of that obtained by OR-AutoTuner. We hypothesize this is from the ECP values entering the agent’s context window, causing it to avoid actions that would worsen ECP.
Figure 3: Optimization trajectory of ASAP7 IBEX (co-optimization, 600 iterations, 12 parameters with tool use). Wirelength (WL), effective clock period (ECP), instance count (Count) are normalized to baseline values from default ORFS settings.
IV-B Multi-objective Optimization
In the case of multiple losses—e.g., $\mathbf{WL}$ and $\mathbf{ECP}$—we instruct the LLM to consider the loss:
$\mathbf{\frac{WL}{WL_{\alpha}}}+\mathbf{\frac{ECP}{ECP_{\alpha}}}$ (1)
with the same substitution, if applicable, for $\mathbf{ECP^{\prime},WL^{\prime}}$ if any of these are unavailable, with the same phrasing adopted above. The experimental results are shown in Table II. Key results are as follows:
  • Sustained gains. Unlike single-objective optimization, where gains observed exceed $13\%$, gains on multi-objective cases are roughly $11.3\%$ on ECP and $10.5\%$ on wirelength simultaneously. This is measured relative to the baselines discussed in subsection IV-C, having $12$ parameters to optimize for $600$ iterations with tool use. Out of $12$ comparisons to OR-AutoTuner, ORFS-agent is superior in $6$ of them, thereby exactly drawing.
  • No worsening. As with single-objective optimization, ORFS-agent does not worsen any metric - either ECP or wirelength, relative to baselines. In fact, both QoRs improve simultaneously.
We visualize the optimization trajectory of instance count alongside wirelength and ECP in Figure 3. We notice that instance count can actually improve for free. To better understand gains on auxiliary metrics (metrics that do not directly enter the optimization function), we examine correlational properties of metrics in Figure 4. We look at how other metrics vary in an optimization run as we carry out ORFS-level runs for ASAP7 IBEX under different parameters and differing optimization goals - namely wirelength, ECP, or a combination of the two as defined above. Some metrics do not “come for free” - ECP and power, for instance, have an inverse relationship with each other.
Figure 4: Correlation of metrics for ASAP7 IBEX: wirelength (WL), effective clock period (ECP), instance area (Area), instance count (Instances) and power (Power). Pearson correlation captures linear relationships; Kendall correlation captures rank-based relationships.
IV-C Baseline comparisons
We obtain baseline solution metrics using the default ORFS flow parameters seen at the cited commit hash. These metrics are shown in Tables III and  IV. Compared to these baselines, ORFS-agent achieves an improvement of over 13% versus the default ORFS flow parameters, when averaged across all six testcases and single-metric (i.e., ECP and wirelength) comparisons. (For all metrics, lower is better.) For additional context, we also provide in Table V the post-route metrics obtained by a proprietary, closed-source tool on ASAP7 circuits; these are quoted from [21]. [^7] Via correspondence with authors of [21], we understand that OpenROAD uses less memory and vCPU resource than the commercial tool, by factors of 5$\times$ or more.
CTSWL CTSECP DRWL ECP
Sky130HD-IBEX 550963 10.84 808423 11.54
ASAP7-IBEX 93005 1308 115285 1361
Sky130HD-AES 428916 5.34 589825 4.72
ASAP7-AES 61103 432 75438 460
Sky130HD-JPEG 1199090 8.00 1374966 7.73
ASAP7-JPEG 266510 1096 300326 1148
TABLE III: Parameter values for original circuits, no changes.
Instance area Count Power PDP
Sky130HD-IBEX 192784 20944 0.097 1.12
ASAP7-IBEX 2729 21831 0.057 77.58
Sky130HD-AES 122361 18324 0.411 1.94
ASAP7-AES 2046 17693 0.149 68.54
Sky130HD-JPEG 541327 65670 0.811 6.27
ASAP7-JPEG 7904 68287 0.138 158.42
TABLE IV: Derived values for original circuits.
Design Tool ECP Area Power
AES Comm 582 40679 0.3498
ORFS* 697 58995 0.2361
IBEX Comm 1123 68843 0.7949
ORFS* 1654 47969 0.4739
JPEG Comm 616 156700 0.7522
ORFS* 899 182624 0.7944
TABLE V: Results for AES, IBEX, and JPEG on optimized OpenROAD (ORFS*) and proprietary tools (Comm) as cited from [21].
IV-D Constrained Optimization under Natural Language Commands
A human engineer is typically given commands significantly different from the utility functions used by OR-AutoTuner. Specifically, a common class of optimization commands relayed to humans takes the form: “Optimize metric $X$ as much as possible, while not allowing metrics $A,B,C$ to deteriorate by more than $2\%$.” This percentage value is with reference to the previously mentioned baseline values in Tables III and  IV. This form of constrained optimization is generally not straightforward with existing OR-AutoTuner setups. LLMs can attempt such optimization by incorporating these constraints directly into their context. We study this phenomenon in Table VI, where we optimize $ECP$ and $WL$ subject to a maximum worsening of $2\%$ and $4\%$ in other metrics of interest—Area, Instance Count, Power, and PDP. We term the maximum worsening allowed as “leeway”. We observe that this often improves multiple metrics simultaneously. This avoids the problem of, e.g., worsening power values that we observed in Fig 4. Comparisons to OR-AutoTuner and further discussion appears in the appendix, alongside tests for robustness and prompt-level sensitivity. ORFS-agent respects the constraints (in terms of maximum $2\%$ or $4\%$ worsening) imposed. Simultaneously, it improves the metrics to be optimized relative to baseline values provided previously.
4 parameters and no tools
Metric WL Optimization ECP Optimization
2% leeway 4% leeway 2% leeway 4% leeway
WL -8.18% -8.51% -7.62% -6.34%
ECP -1.07% 2.51% -2.98% -3.06%
Area -1.36% -1.68% -1.18% -0.80%
Count -1.55% -1.70% -1.35% -1.17%
PDP -2.03% -9.04% -5.12% -0.05%
Power -0.97% -11.27% -2.20% 3.11%
12 parameters with tools
Metric WL Optimization ECP Optimization
2% leeway 4% leeway 2% leeway 4% leeway
WL -11.24% -12.05% -7.9% -8.3%
ECP -3.02% -2.12% -4.18% -4.46%
Area -2.17% -1.92% -1.12% 0.56%
Count -1.62% -1.74% -1.31% -1.06%
PDP -3.25% -0.27% -6.88% -0.77%
Power -0.24% 1.87% -2.81% 3.86%
TABLE VI: Constrained optimization results. Red positive numbers indicate worsening, i.e., an increase relative to baseline, and the “leeway” indicates the maximum worsening allowed.
V Conclusion and Future Directions
In this work, we propose ORFS-agent, a plug-in open-source framework that integrates LLMs into OpenROAD for hyperparameter optimization. Despite using 40% fewer iterations, ORFS-agent achieves results comparable to OR-AutoTuner. ORFS-agent adopts a modular, provider-agnostic framework, which makes it user-friendly and easy to adopt. ORFS-agent can switch seamlessly among LLM providers (e.g., Anthropic, OpenAI, Gemini) without reconfiguration. Additionally, ORFS-agent avoids the cost and complexity of fine-tuning and deployment, and requires no retraining or infrastructure changes when a new LLM is released. Thus, our ORFS-agent can instantly benefit from model breakthroughs. Unlike fine-tuning-based approaches, ORFS-agent can rapidly and cost-effectively adopt stronger models, making it both competitive and appealing.
Our future work will explore the following directions. (i) The field has seen the emergence of reasoning-based models, which use a longer chain of thought process before making decisions. These reasoning-based models are then fine-tuned to task-specific agentic models. We aim to integrate such agentic models into ORFS-agent to enhance its decision-making depth and adaptability. (ii) Our current framework uses a fixed set of pre-defined tools. In future work, we will enable the LLM agent to generate its own tools dynamically, which avoids the hand-coding of specific tools and increases the flexibility in handling different circuits and technology nodes. (iii) We plan to integrate ORFS-agent with search APIs and RAG tools, which enable the agent to access documentation, literature or previous published results. And (iv) we will evaluate ORFS-agent on larger-scale circuits and with more hyperparameters to assess its scalability and robustness. In combination with open-sourcing and OpenROAD integration, we believe that this work will add to the foundations for new research on LLMs for physical design and EDA.
Appendix A EDA Flow and OpenROAD
The EDA flow is highly iterative , and suboptimal outcomes at one step may require going back to earlier steps for re-optimization (e.g., re-synthesizing netlist or adjusting placement). Each step generates logs and metrics that can be integrated into ML pipelines to guide subsequent decisions.
Alongside commercial tools, the OpenROAD project [3] provides a fully open-source, end-to-end RTL-to-GDS solution. It integrates logic synthesis, floorplanning, placement, clock tree synthesis and routing into a fully automated framework. OpenROAD is particularly well-suited for ML research due to the following advantages:
  • Transparency and Instrumentation: Researchers have direct access to internal data structures and logs, enabling detailed data collection and analysis.
  • Reproducibility: The open-source nature of OpenROAD facilitates replication by third parties.
  • Extensibility: APIs are available for integrating custom ML modules. For instance, OR-AutoTuner [22] applies Bayesian optimization (BO) to tune synthesis and place-and-route (SP&R) recipes.
OpenROAD provides a powerful sandbox for exploring data-driven approaches through the digital implementation process. Its open-source foundation enables researchers to modify the details of each implementation step and apply advanced ML techniques for parameter tuning, downstream prediction and flow optimization. In contrast, commercial tools are nearly unusable for open-source academic ML research, due to extremely high license fees and legal constraints on sharing even partial outputs.
Problems with Working with Proprietary Tools. Commercial EDA tools from industry leaders like Synopsys and Cadence are integral to modern chip design, offering comprehensive solutions that cover the entire design flow, including logic synthesis, simulation, and verification. These tools are renowned for their robustness and efficiency, making them the go-to choice for many semiconductor companies.
However, for machine learning (ML) practitioners and researchers, these proprietary tools present several challenges:
  • Closed Architectures: The proprietary nature of these tools means their internal workings are not accessible. This opacity hinders the ability to customize or extend functionalities, which is often necessary for integrating novel ML algorithms or for conducting in-depth research that requires toolchain modifications. Indeed, even the partial outputs cannot be shared without violating legal constraints, a strong contrast from the open-source ecosystem of Machine Learning.
  • High Costs: Licensing fees for commercial EDA tools can be prohibitively expensive, especially for academic institutions, startups, or individual researchers operating under budget constraints. This financial barrier limits the accessibility of these tools for experimental and exploratory ML applications.
  • Limited Flexibility: The closed file structures and lack of transparency in proprietary tools make it difficult to add customized features or promptly address bugs. This rigidity can impede the iterative and experimental nature of ML research, where rapid prototyping and testing of new ideas are crucial.
These constraints underscore the need for open-source alternatives that provide the flexibility, transparency, and cost-effectiveness required for ML-driven EDA research and development.
Open-Source EDA and OpenROAD as an ML Testbed. To systematically study LLM- and ML-driven EDA optimizations, we focus on the OpenROAD project [3]. In contrast to commercial EDA offerings, OpenROAD is fully open-source, thus avoiding copyright restrictions and enabling access to the internal details of its place-and-route algorithms, log files, and data structures. This open access allows the research community to replicate and extend experiments, train and evaluate new models, and publish results without encountering proprietary barriers. As a result, OpenROAD constitutes a natural choice for academic and industrial collaboration, serving as a reproducible testbed in which machine learning and LLM-based agents can be integrated. The availability of its source code and APIs facilitates seamless instrumentation for collecting training data, real-time metrics, and debugging information—all of which are vital for applying advanced ML or agent-based strategies.
In short, leveraging OpenROAD allows us to (1) faithfully represent the full digital implementation flow (floorplanning, placement, clock tree synthesis, routing, etc.), (2) integrate with a community-driven research ecosystem, and (3) avoid opaque “black-box” workflows that impede effective ML experimentation. This environment provides an ideal launchpad for exploring the potential of LLM-based EDA agents that can dynamically interpret intermediate results and fine-tune the flow for improved outcomes.
Appendix B Additional Experimental Details
B-A Details on Parameters, Metrics Tracked, and Surrogates
The details of parameters are shown in Table VII. Table VIII summarizes metrics tracked and surrogates.
Parameter Description Type Range
Clock Period Target clock period (ns/ps) Float (0, $\infty$)
Core Utilization % core utilization Integer [20, 99]
TNS End Percent % violating endpoints to fix Integer [0, 100]
Global Padding Global placement padding level Integer [0, 3]
Detail Padding Detailed placement padding level Integer [0, 3]
Enable DPO Detailed placement optimization Binary {0, 1}
Pin Layer Adjust Routing adjust for metal2/3 Float [0.2, 0.7]
Above Layer Adjust Routing adjust for metal4 and above Float [0.2, 0.7]
Density Margin Add-On Global density margin increase Float [0.00, 0.99]
Flatten Hierarchy Flatten design hierarchy Binary {0, 1}
CTS Cluster Size Number of sinks per CTS cluster Integer [10, 40]
CTS Cluster Diameter Physical span of each CTS cluster Integer [80, 120]
TABLE VII: Overview of optimization parameters for both 4- and 12-parameter tuning. The top four rows in bold form the 4-parameter case.
Optimized Metrics
ECP Final stage effective clock period, defined as the clock period minus the worst slack at the final stage of the build process.
DRWL ($WL$) Post-DR wirelength.
Surrogates of Optimized Metrics
CTSECP ($ECP^{\prime}$) Effective clock period calculated at CTS
CTSWL ($WL^{\prime}$) Estimated wirelength at CTS
Other Desirable Figures of Merit
Instance Area Total area of instances at the end of the optimization process
Instance Count Number of instances at the end of the optimization process
EPC (PDP) Energy per clock cycle, a.k.a. Power Delay Product (PDP)
Power Power consumption at the end of the optimization process
TABLE VIII: Definitions of key metrics.
The Clock Period is specified to one decimal place. Core Utilization and TNS End Percent are integers between 00 and $100$ inclusive, while LB Add On Place Density (henceforth LB Add On) is a real number specified to two decimal places between 00 and $1$. Thus, the latter three variables have $100$ discrete values each, while the Clock Period remains unbounded.
The Effective Clock Period is defined as the clock period minus the worst slack at the finish stage. These metrics, along with their CTS-level surrogates and other desirable figures of merit, are defined in Table VIII. These are based on the METRICS2.1 suite. The Optimized metrics directly enter the natural language optimization process for the LLM. The Surrogates are included only if the primary metrics are unavailable due to timeouts. The remaining parameters—Instance Area, Instance Count, PDP, Power—are not directly optimized but are tracked as meaningful quantities. In the expanded separate configuration (“12-variable setting”), we introduce eight additional tunable parameters that guide the physical design flow more comprehensively. Specifically:
  • GP Pad: Cell padding added during global placement to ensure sufficient spacing.
  • DP Pad: Refinement of cell spacing in detailed placement for higher-quality placement.
  • DPO Enable: A binary toggle for detailed placement optimization.
  • Pin Layer Adjust: Routing resource adjustment on metal2 and metal3 layers.
  • Above Layer Adjust: Similar routing resource adjustment for metal4 and higher layers.
  • Flatten: A binary toggle that flattens design hierarchy.
  • CTS Cluster Size: Target number of sinks grouped in each clock tree cluster.
  • CTS Cluster Diameter: Target physical span of each clock tree cluster.
Appendix C Analysis and Ablations under ASAP7 IBEX
Effects of Iterations and Indirect Optimization. We will consider the results of the joint $WL,ECP$ optimization as a function of the total serial iterations (keeping the number of parallel runs fixed at $25$). The results are shown in Table IX, along with the other relevant figures of merit from Table VIII. Note that each row represents a separate run that is run upto the corresponding number of serial iterations with $25$ parallel runs. That is, $5$ should be considered as a total of $25\times 5=125$ OpenROAD calls.
WL ECP Area Count PDP Power
5 108267 1353 2711 21614 78.474 0.058
10 108530 1334 2701 21529 73.37 0.055
15 111557 1294 2742 21875 89.286 0.069
20 110776 1293 2735 21831 89.217 0.069
25 112495 1298 2729 21796 85.734 0.066
30 108409 1328 2711 21614 77.024 0.058
TABLE IX: Optimization trajectory $5-30$ iterations, no tool use, 4 parameters.
This phenomenon can be well understood in the light of the strong inter-correlations of the data, presented in the main text.
Reversibility. From Table IX, it is clear that optimizing wirelength and Effective Clock Period (ECP) jointly also optimizes other parameters. We now reverse this procedure to analyze the effect on wirelength and ECP when any one of these parameters is optimized individually. The results of this process are presented in Table X. Notably, this process appears to be asymmetric—these iterations yield noticeably worse outcomes on ECP while resulting in significantly lower power consumption.
Count Area PDP Power WL ECP
Count 21419 2679 68.632 0.046 109084 1492
Area 21437 2674 70.952 0.049 117411 1448
PDP 21432 2686 68.310 0.046 106070 1485
Power 21516 2689 68.540 0.046 107350 1490
WL 21540 2701 76.828 0.058 106044 1324
ECP 21799 2748 86.483 0.067 113854 1281
TABLE X: Results of optimizing the row-level objective on column-level outcomes, ASAP7 IBEX, no tool use, 4 parameter setting.
To ensure that our results stay consistent across tool use choices, we replicate this experiment on ASAP7 IBEX with 12 parameters and tool use enabled in Table XI.
Count Area PDP Power WL ECP
Count 21356 2683 75.28 0.054 112984 1394
Area 21382 2642 76.21 0.058 120811 1314
PDP 21486 2706 59.41 0.047 107822 1264
Power 21416 2683 59.06 0.046 104622 1284
WL 21682 2672 99.68 0.078 97305 1278
ECP 21674 2684 102.34 0.082 105562 1248
TABLE XI: Results of optimizing the row-level objective on column-level outcomes, ASAP7 IBEX, tool use and 12 parameter setting.
Constrained Optimization under Natural Language Commands. We presented the results of the constrained optimization setup in the main text. Here, we continue on this vein. A human engineer is typically given commands significantly different from the utility functions used by OR-AutoTuner. Specifically, a common class of optimization commands relayed to humans takes the form:
“Optimize metric $X$ as much as possible, while not allowing metrics $A,B,C$ to deteriorate by more than $2\%$.”
This form of constrained optimization is generally not straightforward with existing OR-AutoTuner setups. LLMs can attempt such optimization by incorporating these constraints directly into their context. We study this phenomenon in Table XII and compare it to OR-AutoTuner in Tables XIII and  XIV, where we optimize $ECP$ and $WL$ subject to a maximum worsening of $2\%$ and $4\%$ in other metrics of interest—Area, Instance Count, Power, and PDP. These percentage values are always with respect to the baseline values generated from default OR-AutoTuner parameters.
4 parameters and no tools
Metric WL Optimization ECP Optimization
2% leeway 4% leeway 2% leeway 4% leeway
WL -8.18% -8.51% -7.62% -6.34%
ECP -1.07% 2.51% -2.98% -3.06%
Area -1.36% -1.68% -1.18% -0.80%
Count -1.55% -1.70% -1.35% -1.17%
PDP -2.03% -9.04% -5.12% -0.05%
Power -0.97% -11.27% -2.20% 3.11%
12 parameters with tools
Metric WL Optimization ECP Optimization
2% leeway 4% leeway 2% leeway 4% leeway
WL -11.24% -12.05% -7.9% -8.3%
ECP -3.02% -2.12% -4.18% -4.46%
Area -2.17% -1.92% -1.12% 0.56%
Count -1.62% -1.74% -1.31% -1.06%
PDP -3.25% -0.27% -6.88% -0.77%
Power -0.24% 1.87% -2.81% 3.86%
TABLE XII: Constrained optimization results. Red positive numbers indicate worsening, i.e. an increase relative to baseline, and the “leeway” indicates the maximum worsening allowed.
Metric WL Optimization ECP Optimization
2% leeway 4% leeway 2% leeway 4% leeway
WL -9.03% -5.69% -3.23% -1.41%
ECP 1.40% 1.03% -1.31% -2.60%
Area -1.59% -1.22% -0.39% -0.08%
Count -1.80% -1.71% 0.58% -0.02%
PDP -4.16% -7.87% -0.55% -0.06%
Power -5.49% -8.81% 0.77% 2.61%
TABLE XIII: Constrained optimization results for OR-AutoTuner (with 375 iterations, four parameters). Red positive numbers indicate worsening, i.e., an increase relative to baseline, and the “leeway” indicates the maximum worsening allowed.
Metric WL Optimization ECP Optimization
2% leeway 4% leeway 2% leeway 4% leeway
WL -12.53% -9.26% 1.59% -7.04%
ECP 1.07% 2.46% -0.46% -2.67%
Area -1.86% -1.65% -0.49% -1.95%
Count -2.0% -1.47% 0.16% -1.57%
PDP -7.17% -4.17% -1.10% -2.31%
Power -8.15% -6.47% -0.65% 0.37%
TABLE XIV: Constrained optimization results for OR-AutoTuner (with 1000 iterations, 12 parameters). Red positive numbers indicate worsening, i.e., an increase relative to baseline, and the “leeway” indicates the maximum worsening allowed.
The $4\%$ tolerance case naturally gives rise to a Pareto frontier scatterplot of Effective Clock Period (ECP) versus wirelength (WL), which is shown in Figure 5.
Figure 5: Pareto frontier of ECP vs. WL under $4\%$ tolerance, relative to the baseline. The blue line represents the case with 4 parameters and no tool use, and the red line represents the case with 12 parameters and tool use.
Appendix D Ablations for Data Leakage, Confidence and Robustness
We begin by analyzing statistical significance in table XV, and testing prompt-level changes in tables XVI and  XVII. The “role” here should be understood as the role of an EDA expert that the LLM is asked to pose as. The details exist separately on the level of PDK / platform, circuit, and parameters. Each of these is ablated against. The reader should carefully also check the implementation codebase provided with the paper to understand the nature of this metadata.
TABLE XV: Statistical significance over 20 independent ORFS-agent runs on the ASAP7 IBEX circuit. Percentile columns report the $k$-th worst value ($k\!=\!1,5,10,15,19$) among the 20 outcomes; “Reported” is the single-run value quoted in the main paper.
Metric $5$-th $25$-th Median $75$-th $95$-th Reported
WL 99820 98512 97360 96075 94980 97305
ECP* 1335 1301 1275 1251 1223 1278
ECP 1294 1271 1246 1224 1197 1248
WL* 108430 107180 105598 104125 102470 105562
WL** 105230 103720 102178 10162 99740 102205
ECP** 1311 1283 1259 1232 1201 1260
TABLE XVI: Ablation on domain-knowledge content in prompts for ASAP7 IBEX. “PDK” removes technology-node details; “circuit” removes design-specific details; “Both” drops both. Higher WL/ECP indicate worse quality of result.
Configuration WL ECP* ECP WL* WL** ECP**
Baseline 97305 1278 1248 105562 102205 1260
No PDK 99874 1385 1296 107284 105210 1275
No circuit 98743 1368 1301 106938 104782 1280
Both removed 106044 1324 1281 113854 114576 1287
TABLE XVII: Sensitivity of ORFS-agent to prompt simplifications on ASAP7 IBEX. “Removed role” deletes the system-role description; “Removed param” deletes parameter-guidance text.
Prompt variant WL ECP* ECP WL* WL** ECP**
Baseline 97305 1278 1248 105562 102205 1260
No role 98044 1292 1255 106114 102934 1267
Blank params 99130 1310 1270 107039 103848 1278
Throughout the process, the large amount of data from Anthropic raises concerns about whether the LLM is simply caching or memorizing what it observes. This includes the possibility of interactions with other OpenROAD users who may be testing Claude models on similar outputs. We test these concerns in the following ways:
Circuit Obfuscation. To assess potential memorization, we obfuscate the name of the ASAP7-IBEX circuit and subject it to the same iteration vs. performance ablation test. The circuit name is changed to obfs, while the design platform input to the ORFS-agent remains ASAP7. Additionally, configuration settings are modified to rename the top-level module from “IBEX” to obfs. The ORFS-agent is also restricted from grepping log outputs that refer to the original name or netlist identifiers. The results appear in Table XVIII and Table XIX.
WL ECP Area Count PDP Power
5 110605 1311 2731 21728 86.526 0.066
10 113186 1304 2741 21972 89.976 0.069
15 110358 1296 2735 21735 88.128 0.068
20 111918 1281 2746 21854 89.670 0.070
25 110965 1283 2733 21732 85.961 0.067
30 112645 1273 2746 21882 90.383 0.071
TABLE XVIII: Obfuscated results—ASAP7 IBEX, no tools and with 4 parameters.
WL ECP Area Count PDP Power
5 114861 1316 2716 22015 84.22 0.064
10 105788 1294 2728 21986 90.58 0.070
15 106724 1286 2735 21924 87.45 0.068
20 106245 1278 2704 21864 84.35 0.066
25 103871 1276 2694 21882 90.60 0.071
30 103244 1272 2711 218008 94.13 0.074
TABLE XIX: Obfuscated results—ASAP7 IBEX, with tools and with 12 parameters, including PDP = ECP $\times$ Power.
Statistical Significance. We run the ASAP7-IBEX optimization for $ECP,WL$ $10$ times, with results in Table XX and  XXI. Even the worst of the $10$ produces a result that is superior in terms of ECP against OR-AutoTuner. Each individual row indicates the outcome with the lowest ECP in the final $25$ iterations of an individual run.
WL ECP Area Count PDP Power
Run 1 116479 1290 2744 21781 90.300 0.070
Run 2 110232 1286 2736 21775 90.020 0.070
Run 3 121092 1287 2778 22181 91.377 0.071
Run 4 114576 1287 2746 22013 84.942 0.066
Run 5 109107 1303 2720 21636 80.786 0.062
Run 6 111054 1275 2740 21816 91.800 0.072
Run 7 116722 1279 2758 22098 89.530 0.070
Run 8 115933 1281 2766 21918 90.951 0.071
Run 9 109230 1309 2724 21656 82.467 0.063
Run 10 109342 1299 2728 21734 88.332 0.068
TABLE XX: $10$-fold runs, ASAP7 IBEX, 4 parameters and no tools.
WL ECP Area Count PDP Power
Run 1 102289 1288 2726 21688 100.464 0.078
Run 2 102677 1285 2715 21722 102.800 0.080
Run 3 104200 1254 2705 21782 102.828 0.082
Run 4 103384 1270 2674 22092 109.220 0.086
Run 5 104922 1268 2678 21954 84.956 0.067
Run 6 102874 1269 2691 21894 91.368 0.072
Run 7 103982 1289 2712 21866 95.386 0.074
Run 8 103581 1275 2732 21895 96.900 0.076
Run 9 106742 1248 2725 21906 88.608 0.071
Run 10 105864 1252 2698 21967 87.640 0.070
TABLE XXI: 10-fold runs, ASAP7 IBEX, 12 parameters and with tools.
Robustness Testing. We perform two layers of robustness testing. First, we add noise to every measurement: we multiply every observation $O$, at random, with the variable $1+X_{O}$, where each $X_{O}$ is an i.i.d. uniform random variable from $[-0.05,0.05]$. Next, we consider the case where our specification of the surrogate has gone astray - it is a common occurrence in EDA for the variable being optimized to not really track what is desired. For this, we change all the CTS-level observations to be multiplied by $1+X_{CTS}$, where $X_{CTS}$ is an uniform random variable i.i.d drawn from $[-0.2,0.2]$. The results on ASAP7 IBEX are presented below in Table XXII.
Run WL CP Area Count PDP Power
Overall perturbation, per objective case
WL 104427 1318 2741 21782 92.26 0.070
ECP 106872 1268 2708 21905 98.90 0.078
WL & ECP6 107848 1274 2685 21802 96.82 0.076
CTS perturbation, per objective case
WL 106898 1305 2673 21917 93.96 0.072
ECP 112824 1272 2711 21904 96.67 0.076
WL & ECP 109873 1281 2705 21877 93.51 0.073
TABLE XXII: Perturbed runs, 12 parameters, tool use permitted.
Appendix E Global Context, Initialization, and the Toolbox
The FunctionalORFS-agent maintains a GLOBALCONTEXT containing the design environment details (PDK, circuit, task specifics, etc.). Each LLM call in the pipeline includes this global context, plus the relevant logs and data extracted at runtime. The GLOBALCONTEXT is initialized as the prompt, defined below.
LLM calls made within each part of an iteration - i.e., INSPECT, etc. - add the LOCALCONTEXT of that particular iteration. However, they are not necessarily passed on to the GLOBALCONTEXT. As we move from iteration $j$ to $j+1$, GLOBALCONTEXT is modified to only include a small subset of all the LOCALCONTEXT within the particular iteration $j$.
E-A Initialization
At the very first iteration, the prompt is initialized based on a set of the following inputs:
  • The design platform (e.g., ASAP7).
  • The circuit under consideration (e.g., AES).
  • The optimization task (e.g., minimize wire length (WL)).
  • The inputs, which include design parameters (e.g., the 12 parameters such as core utilization, etc.).
  • The outputs, specifying the variable sets (e.g., detailed WL) and variable sets of its surrogate(s) (e.g., CTS WL).
  • The exact output to optimize from the variables (e.g., -WL for maximization or WL for minimization).
  • The domains of the inputs, including whether they are Boolean, integers with min/max bounds, or continuous values converted to integers (e.g., converting float ranges like 0.01–0.99 to integer ranges 1–99, and scaling back to original values for PARAMETERS generation).
  • Suggested ranges for inputs.
E-B Toolbox
We define three separate kinds of tools. All of them require a Description, a set of ideally typed arguments, and a return object Return with description and type specifications to be efficiently made a part of LLM function-calling frameworks. All of them are summarized in section F.
INSPECT Tools. These tools enable quick analysis of numeric datasets by computing summary statistics, correlations, outlier indices, or performing principal component analysis. They do not consume raw data token-by-token but instead return concise analytical results that aid downstream decision-making.
OPTIMIZE Tools. These tools propose new points (parameter sets) to explore in an optimization process. Each one uses a different strategy—ranging from random sampling and grid-based approaches to more sophisticated algorithms like Bayesian optimization or genetic algorithms—to balance exploration and exploitation.
AGGLOM Tools. These tools agglomerate (reduce or cluster) a large set of candidate solutions into a smaller or more representative subset. Methods include selecting the Pareto front for multi-objective problems, maximizing coverage or diversity, or clustering to find the most central or diverse candidates.
Appendix F Tool Details
INSPECT Tools
AnalyzeManifold analyzes the underlying manifold structure using PCA, TSNE, and MDS. It takes X: np.ndarray and an optional config: Dict, and returns a summary as a Dict.
AnalyzeLocal examines local structure using LOF and DBSCAN. It accepts X: np.ndarray and an optional config: Dict, and returns a summary Dict.
InspectDistribution analyzes statistical properties of input/output data. It requires X: np.ndarray, Y: np.ndarray, and optionally Y_surrogate: np.ndarray. The output is a summary: Dict.
InspectStructure inspects structural properties and gives model recommendations. Inputs include X: np.ndarray, Y: np.ndarray, and optionally config: Dict. It returns a summary: Dict.
OPTIMIZE Tools
CreateModel creates and configures a Gaussian Process model. It takes X: np.ndarray, y: np.ndarray, noise_level: float, and kernel_type: str, returning a model of type GaussianProcessRegressor.
CreateKernel generates a kernel from a given specification. The required inputs are kernel_spec: str and input_dim: int. It outputs a GPR kernel.
ExpectedImprovement calculates the Expected Improvement (EI) acquisition function. Inputs include mu: np.ndarray, std: np.ndarray, and y_best: float. The return is ei: np.ndarray.
HandleSurrogate processes and combines true and surrogate data. It uses X: np.ndarray, y: np.ndarray, and surrogate_values: np.ndarray, and returns a tuple (X, y_combined, uncertainty).
LatinHypercube generates Latin Hypercube samples. The inputs are n_points: int and n_dims: int, returning samples: np.ndarray.
AGGLOM Tools
SelectPoints selects points based on a specified method. It accepts X: np.ndarray, quality_scores: np.ndarray, method: str, and n_points: int, and returns selected_indices: np.ndarray.
HybridSelect combines quality and diversity for point selection. Required arguments are X: np.ndarray, quality_scores: np.ndarray, distance_matrix: np.ndarray, and n_points: int, with return selected_indices: np.ndarray.
EntropySelect uses entropy-based diversity for selection. It takes X: np.ndarray, quality_scores: np.ndarray, and n_points: int, returning selected_indices: np.ndarray.
GraphSelect applies graph-based diversity metrics for point selection. It accepts X: np.ndarray, quality_scores: np.ndarray, and n_points: int, returning selected_indices: np.ndarray.
CreateQualityScores computes quality scores from model predictions. Inputs include X: np.ndarray, y: np.ndarray, model_predictions: np.ndarray, and optionally model_uncertainties: np.ndarray. It returns quality_scores: np.ndarray.
Each of the tools used here can be implemented using standard python packages. For instance, scikit-learn, numpy and related Python packages offer support for Gaussian Processes, t-SNE modeling, and so on. By combining these existing libraries with domain-specific logic, we create a modular, maintainable toolbox that reflects the iterative exploration process a human data scientist would perform —streamlining tasks like outlier detection, dimensionality reduction, and optimization-driven design.